1. calculate the volume O2 (g) measured at 20oC and 120.0 KPa that can be obtained from the decomposition of 15.0 g of KNO3 from the reaction: 2 KNO3 (s) --> 2 KNO2 (s) + O2 (g)
  2. How many liters of O2 (g) measured at 30oC and 110.0 KPa is needed to completely burn 20.0 liters of butane, C4H10(g) + 13/2 O2 (g) -> CO2(g) + 5 H2O(/).
1
Expert's answer
2021-01-05T06:02:24-0500

1.

number of moles in 15g of KNO3 = 15/101 = 0.149 moles


"2 KNO_{3 (s)} \to 2 KNO_{2 (s)} + O_{2 (g)}"

At STP, 2 moles of KNO3 decompose to give 1 mole of O2

This means that 0.149 moles of KNO3 will decompose to give 0.0745 ("\frac{0.149}{2}") moles of O2


T = 20°C = 293K

P = 120.0kPa = 120,000 Pa

n = 0.0745 moles


using the ideal gas law,

PV = nRT


"V = \dfrac{nRT}{P} = \dfrac{0.0745 ×8.314× 293}{120000}"


= 0.00151m³ = 1.51dm³ (1.51L)



2.

20.0 litres of butane = 20/22.4 moles = 0.893 moles


"C_4H_{10(g)} + \frac{13}2 O_{2 (g)} \to CO_{2(g)} + 5 H_2O_{(l)}"

According to the reaction above,

1 mole of C4H10 uses 6.5 moles of O2 to combust completely

This means that 0.893 moles of C4H10 uses 5.8 (0.893×6.5) moles of O2 to combust completely


T = 30°C = 303K

P = 110.0kPa = 110,000 Pa

n = 0.893 moles


using the ideal gas law,

PV = nRT


"V = \dfrac{nRT}{P} = \dfrac{5.8 ×8.314× 303}{110000}"


= 0.1328m³ = 132.8dm³ = 132.8L